OpenStudy (anonymous):

using 15degrees=45degrees-30 degrees what would be the sin, cos and tan of 15 degrees?

OpenStudy (amistre64):

sin(a+ -b) = sin(a)cos(-b) + sin(-b)cos(a) cos(a+ -b) = cos(a)cos(-b)-sin(a)sin(-b) and tan = sin/cos

OpenStudy (amistre64):

tan(a+b) tan(a)tan(b) ---------- 1+tan(a)tan(b) maybe

OpenStudy (mertsj):

sin(45-30)=sin45cos30-cos45sin30 cos(45-30)=cos45cos30+sin45sin30

OpenStudy (anonymous):

whats a and b?

OpenStudy (mertsj):

a and b are representative of whatever angles you are working with. In this case a is 45 and b is 30

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

well, i went ahead and used -b for -30

OpenStudy (amistre64):

otherwise Im sure I would have messed up the operation in the middle :)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

im having problems with tangent